# Distributed Learning Dynamics Converging to the Core of -Matchings

### v1

@@ Abstract
-matching is a special case of matching problems where nodes can join multiple matchings with the degree of each node constrained by an upper bound, the nodeβs -value. The core solution of a bipartite -matching is both a matching between the agents respecting the upper bound constraint and an allocation of the value of the edge among its nodes such that no group of agents can deviate and collectively gain higher allocation. We present two learning dynamics that converge to the core of the bipartite -matching problems. The first dynamics are centralized dynamics in the nature of the Hungarian method, which converge to the core in a polynomial time. The second dynamics are distributed dynamics, which converge to the core with probability one. For the distributed dynamics, a node maintains only a state consisting of (i) its aspiration levels for all of its possible matches and (ii) the matches, if any, to which it belongs. The node does not keep track of its history nor is it aware of the environment state. In each stage, a randomly activated node proposes to form a new match and changes its aspiration based on the success or failure of its proposal. At this stage, the proposing node inquires about the aspiration of the agent it wants to match with to calculate the feasibility of the match. The environment matching structure changes whenever a proposal succeeds. A state is absorbing for the distributed dynamics if and only if it is in the core of the -matching.
@@ I INTRODUCTION
Matchings are used widely in multi-agent systems to model setups with pairwise interactions. Edges between agents can describe communication, contracts or assignments. Matchings have been studied in various forms, connection-wise: one-to-one, many-to-one, and many-to-many, and graph-wise: general and bipartite. Constraints and parameters can also be added to limit the number of connections and the weight/cost of these connections. The different forms of matchings allow us to reduce many applications to matching problems. For example, multi-agent task allocation problems can usually be modeled as one form of matching according to the objectives of the application[1]-[3]. Matching applications also include the widely known kidney exchange problem, the medical residents matching program in the United States, and ride-sharing applications [4]-[5].
@@
Various real-life applications can be modeled as many-to-many bipartite matching problems, for instance, online ad placements [6]-[7] and firms and consultants matching [8]. In addition, the authors in [9] modeled data markets where data can be replicated with zero-cost as many-to-many bipartite matching. Furthermore, in the networks field, many problems are related to matching through duality, reduction, or extension, such as edge covering, stable set polyhedron, flow, transportation, and transshipment problems [10].
@@
The author in [11] presented some key considerations for multi-agent systems. Among these considerations are the autonomy of the multi-agent system, the complexity of the decision-making and learning, the communication between agents, the distribution of the agents, and the systemβs security and privacy requirements. The distributed dynamics presented in this paper address a number of these considerations since the dynamics are distributed, require only simple computations from the agents per iteration, and reveal only minimal information about other agentsβ states. The setup we are considering is the bipartite -matching setup, which allows many-to-many matchings with an upper bound on the number of matches per node and at most one match between each pair of nodes. In our dynamics, an agent is randomly activated which proposes to match with another agent if both of their aspirations can be fulfilled by the value of the match. The proposing agent increases or decreases its aspiration according to the success or failure of the proposal and its current matching state. An agent cannot exceed its upper bound of matches, hence if it is already at its capacity for matches, it needs to break off an old match to join a new one.
@@
The presented dynamics converge to the core allocation of -matching problems. The core allocation is an essential solution concept in cooperative games that guarantees that no group of agents can gain more by deviating and forming their own coalition in the game. The core allocation is proven to exist for bipartite -matchings in [12]. We prove that through our distributed dynamics and only pairwise interactions, the dynamics reach a core allocation with probability one.
@@
Our distributed dynamics, the -matching proposals dynamics, extend and build upon the dynamics in [13]-[15]. The dynamics in [13] and [14] address the one-to-one bipartite matching setup and the many-to-one bipartite matching setup, respectively. The dynamics in [15] is applicable to general coalitional games with transferable utility, an example of which is the matching problems on a general graph. Similar work has been carried out in [16] for assignment games and in [17] for superadditive coalitional games. For -matchings, a core allocation can be reached using the dynamics in [15] if it is formulated as a coalitional game, yet it requires the calculations of the maximum -matching value (defined later) for each coalition of agents. Our current proposed dynamics bypass these calculations and use only pairwise interactions to reach the core of the -matching.
@@ II-A Definitions
Set the maximum -matching value of a graph and the associated and functions to be where is the set of -matchings over the graph. The maximum -matching value over a coalition where and is the maximum -matching value of over the subgraph with the associated and functions.
@@
As an illustration, consider a multi-agent task assignment setting, where each agent is able to perform multiple tasks and each task can be performed by several agents. The profit that is realized from an agent performing a task depends on both the agent and the nature of the task. The previous definitions can be interpreted as follows. The two node groups and are the group of agents and the group of tasks, respectively. A connected edge is an agreement that agent performs task The profit gained as agent performs task is the weight of the edge, The profit of matched edges can then be divided among the agents and the tasks. The -value for a task is an upper bound on the maximum number of agents that can perform the task and the -value for an agent is the maximum number of tasks it can perform. The allocation dictates how much gain each agent and task get from the profits collected from all matched edges.
@@
Consider a complete bipartite graph together with a weighing function and a -value Consider and as defined above. Let the degree for all and for each and then is a -copies matching.
@@
A reduction function is a function that maps a pair of copies allocation and -copies matching to a pair of nodes allocation and -matching as follows. If then and Note that is still a -matching as previously defined because for each node
From this paperβs perspective, each node will be expanded to distinct copies, each treated as a node in an expanded graph and gets its own gain from the profit of the matches; a nodeβs gain then is the sum of the gains collected by all of its copies.
@@
Consider the copies-core solution of the previous task assignment setting. For this solution, an allocation is specified for each copy of the node. This finer detailing of the copiesβ allocations shows the gains of each node, robot or task, from each match it is part of, thus, decoupling the individual matchesβ profit distribution. In a copies-core solution, the allocation specified for each copy of a node is only its profit allocation from the match that this copy is a part of.
In the next two sections, we define a centralized polynomial time algorithm as well as distributed dynamics for finding the copies-core solution. Since we are aiming to reach the copies-core solution, we work with the expanded graph as defined above. The following definitions will be used in the subsequent algorithms. Furthermore, we will overload in the dynamics to be the agentsβ allocation aspirations for the copies. In our dynamics, an edge is formed only if these aspirations can be fulfilled by such match.
@@
The following is an assumption that we are going to consider for both the centralized and distributed algorithm.
@@
If a given -matching problem does not satisfy this assumption, then the degree of the violating nodes still will not exceed the number of nodes in the opposite class in any -matching, and hence, the problem remains the same if we decrease the -values of these nodes just enough to satisfy the above assumption.
The algorithms use the aforementioned functions, however, the superscript are suppressed and considered as the matching at the current iteration.
@@ III PATHS TRANSFERS ALGORITHM
Initially, the matching will be empty and the aspirations of all copies will be . Set to be the largest discretization value for all the weights of the edges.
@@
Augmenting path case: If there exists a directed path in from the considered copy to some copy then let and disregarding the directions of the edges. Add edges in to and remove from
This case will lead to decreasing by 1 or 2 depending whether is greater than or equal to
@@
Copies exchange case: If there exists a directed path in from the considered copy to some copy with then let and disregarding the directions of the edges. Add edges in to and remove from
This case will lead to decreasing by 1 as a copy of zero aspiration will be exchanged with a copy of non-zero aspiration in the matched copies.
@@
The inner loop of step must end in a finite time. The inner loop will end if it goes through the augmenting path or copies exchange cases or if the considered copy reaches zero aspiration in the decreasing aspiration case. The inner loop cannot infinitely go through the decreasing aspiration and the aspiration transfer case since the total aspirations of copies in strictly decreases with each implementation of these cases and it is bounded below by zero. Hence, in a finite number of steps, the inner loop will end up in the considered copy reaching zero aspiration and hence decreases or the loop goes through the augmenting path or copies exchange cases.
@@
The outer loop of step must end because strictly decreases with each iteration.
@@
Note that the algorithm can be modified to be polynomial in and and exclude the weight values dependency. To achieve that, the aspiration transfer case should be modified such that the copies aspirations change by the minimum change of aspiration that leads to a new edge in the equality graph to appear or for one of the copies to hit zero aspiration. The decreasing aspiration case should be modified such that the decrease in aspiration be the minimum decrease such that a reconstructed contains an edge from the considered copy to any other copy or its aspiration reaches zero.
@@
In this section, we introduce the -matching proposals dynamics. Intuitively, the algorithm follows these steps
@@
Agents come with arbitrary non-negative initial aspirations for all their copies. The aspirations should lie on a grid of width , a discretization value for all the weights of the edges.
@@
If the current agents are matched, the proposal is skipped, otherwise, the proposing agent asks for and receives, an aspiration value from the other agent.
@@
The receiving agent sends the minimum aspiration of its unmatched copies, if any exist, otherwise, it sends the minimum aspiration of the matched copies. Similarly, the proposing agent considers the minimum aspiration of its unmatched copies, if any exist, otherwise, the minimum aspiration of the matched copies. If the total of the received aspiration in addition to the proposing agentβs considered aspiration raised by is less than or equal to the proposed match weight, a match between the two copies of the considered aspirations is formed, otherwise, the proposal fails.
@@
If a match is successfully formed, the proposing agent increases its aspiration to be the difference between the weight of the edge between the proposer and the receiver and the receiverβs aspiration. If any of the copies in the new match were already matched, it dissolves its previous match to join the new one.
@@
The proof follows three steps illustrated in the following subsections. First, the there exists a positive probability bounded away from zero to reach a copies-core solution from any feasible state within some finite time . Second, the algorithm converges to a copies-core solution with probability one. Third, a state is absorbing if and only if it is a copies-core solution.
I. There exists a positive probability to reach a copies-core solution from any feasible state
@@ II Borel-Cantelli lemma
Proposition 4 implies that there exists time for which the probability of reaching a copies-core solution from any feasible state is at least some that does not depend on the state, following the aforementioned finite sequences. Hence, by the Borel-Cantelli lemma, the algorithm reaches a copies-core solution with a probability one.
III. A state is absorbing if and only if it is a copies-core solution
@@
and the zero-gain for unmatched copies condition is satisfied. Otherwise, such that and due to Assumption there exists a node in the opposite class that is not connected to through any copy. If proposes to match with such node, the proposal should fail since the state is absorbing and the matching structure cannot change, yet one copy of from will decrease its aspiration, hence the state is not absorbing, which provides a contradiction.
@@
Given , then and the stability against pairwise deviation is satisfied. Otherwise, and such that thus if proposes to or vice versa, the proposal will succeed changing the matching and the aspirations. Hence the state is not absorbing, which provides a contradiction.
@@
Assume the state is a copies-core solution at iteration . Assume that at iteration any was chosen to propose to any
@@
If and were already connected: no change occurs.
@@
If and are not connected, then by the stability against pairwise deviation property, and Hence, all proposals will fail and no change can occur in the matching .
In addition, from zero-gain for unmatched copies condition, , then none of the copies can decrease its aspiration.
Similar justification is applied if any proposes to any Hence, is absorbing.
@@ VI ILLUSTRATION
In this section, we demonstrate the distributed learning dynamics on the aforementioned multi-agent task assignment setting. The setup is as follows.
@@
We consider tasks, where
@@
can afford to utilize up to robots
@@
We consider robots, where
@@
each robot has accuracy value where and
@@
can afford to cover up to tasks.
@@
in our setup, we are assuming a negative correlation with the sophistication of the robot and how many tasks it can cover
β
@@
The value of a matching of task with robot is where is proportional with the accuracy of the robot multiplied by the value of the task

### v2

@@ Abstract
-Matching is a special case of matching problems where nodes can join multiple matchings with the degree of each node constrained by an upper bound, the nodeβs -value. The core solution of a bipartite -matching is both a matching between the nodes respecting the upper bound constraint and an allocation of the weights of the edges among the nodes such that no group of nodes can deviate and collectively gain higher allocation. We present two learning dynamics that converge to the core of the bipartite -matching problems. The first dynamics are centralized dynamics in the nature of the Hungarian method, which converge to the core in a polynomial time. The second dynamics are distributed dynamics, which converge to the core with probability one. For the distributed dynamics, a node maintains only a state consisting of (i) the aspiration levels for all of its possible matches and (ii) the matches, if any, to which it belongs. The node does not keep track of its history nor is it aware of the environment state. In each stage, a randomly activated node proposes to form a new match and changes its aspiration based on the success or failure of its proposal. At this stage, the proposing node inquires about the aspiration of the node it wants to match with to calculate the feasibility of the match. The environment matching structure changes whenever a proposal succeeds. A state is absorbing for the distributed dynamics if and only if it is in the core of the -matching.
@@ I INTRODUCTION
Matchings are used widely in multi-agent systems to model setups with pairwise interactions. Edges between agents can describe communication, contracts, or assignments. Matchings have been studied in various forms, connection-wise: one-to-one, many-to-one, and many-to-many, and graph-wise: general and bipartite. Constraints and parameters can also be added to limit the number of connections and the weight/cost of these connections. The different forms of matchings allow us to reduce many applications to matching problems. For example, multi-agent task assignment problems can usually be modeled as one form of matching according to the objectives of the application [1]-[3]. Matching applications also include the widely known kidney exchange problem, the medical residents matching program in the United States, and ride-sharing applications [4]-[5].
@@
Various real-life applications can be modeled as many-to-many bipartite matching problems, for instance, online ad placements [6]-[7] and firms and consultants matching [8]. In addition, the authors in [9] modeled data markets where data can be replicated with zero cost as a many-to-many bipartite matching problem. Furthermore, in the networks field, many problems are related to matching through duality, reduction, or extension, such as edge covering, stable set polyhedron, flow, transportation, and transshipment problems [10].
@@
The author in [11] presented some key considerations for multi-agent systems. Among these considerations are the autonomy of the multi-agent system, the complexity of the decision-making and learning, the communication between agents, the distribution of the agents, and the systemβs security and privacy requirements. The distributed dynamics presented in this paper address a number of these considerations since the dynamics are decentralized, require only simple computations from the agents per iteration, and reveal only minimal information about other agentsβ states. The setup we are considering is the bipartite -matching setup, which allows many-to-many matchings with an upper bound on the number of matches per node and at most one match between each pair of nodes. In our dynamics, an agent is randomly activated which proposes to match with another agent if both of their aspirations can be fulfilled by the value of the match. The proposing agent increases or decreases its aspiration according to the success or failure of the proposal and its current matching state. An agent cannot exceed its upper bound of matches, hence, if it is already at its capacity for matches, it breaks off an old match to join a new one.
@@
The presented dynamics converge to the core allocation of the -matching problems. The core allocation is an essential solution concept in cooperative games that guarantees that no group of agents can gain more by deviating and forming their own coalition in the game. The core allocation is proven to exist for bipartite -matchings in [12]. We prove that through our distributed dynamics and only pairwise interactions, the dynamics reach a core allocation with probability one.
@@
Our distributed dynamics, the -Matching Proposals dynamics, extend and build upon the dynamics in [13]-[15]. The dynamics in [13] and [14] address the one-to-one bipartite matching setup and the many-to-one bipartite matching setup, respectively. The dynamics in [15] apply to general coalitional games with transferable utility, an example of which is the matching problems on a general graph. Similar work has been carried out in [16] for assignment games and in [17] for superadditive coalitional games. For -matchings, a core allocation can be reached using the dynamics in [15] if it is formulated as a coalitional game, yet it requires the calculations of the maximum -matching value (defined later) for each coalition of agents. Our current proposed dynamics bypass these calculations and use only pairwise interactions to reach the core of the -matching.
@@ II-A Definitions
Set the maximum -matching value of a graph and the associated and functions to be where is the set of -matchings over the graph. The maximum -matching value over a coalition where and is the maximum -matching value over the subgraph with the associated and functions.
@@
As an illustration, consider a multi-agent task assignment setting, where each agent is able to perform multiple tasks and each task can be performed by several agents. The profit that is realized from an agent performing a task depends on both the agent and the nature of the task. The previous definitions can be interpreted as follows. The two node groups and are the group of agents and the group of tasks, respectively. A connected edge is an agreement that agent performs task The profit gained as agent performs task is the weight of the edge joining them, The profit of matched edges can then be divided among the agents and the tasks. The -value for a task is an upper bound on the maximum number of agents that can perform the task and the -value for an agent is the maximum number of tasks it can perform. The allocation dictates how much gain each agent and each task get from the profits collected from all matched edges.
@@
Consider a complete bipartite graph together with a weighing function and a -value Consider and as defined earlier. Let the degree of a node for all and for each and then is a -copies matching.
@@
A reduction function maps a pair of copies allocations and -copies matching to a pair of nodes allocation and nodes -matching as follows. If then and Note that is still a -matching as previously defined because for each node
From this paperβs perspective, each node is expanded as distinct copies, each treated as a node in an expanded graph and gets its own gain from the profit of the matches; a nodeβs gain then is the sum of the gains collected by all of its copies.
@@
Consider the copies-core solution of the preceding task assignment setting. The copies-core solution in this setting specifies the allocations of the profit for each copy of the nodes. This finer detailing of the copiesβ allocations shows the gains of each node, robot or task, from each match it is part of, thus, decoupling the individual matchesβ profit distribution. In a copies-core solution, the allocation specified for each copy of a node is only its profit allocation from the match that this copy is a part of.
In the next two sections, we define a centralized polynomial time algorithm as well as distributed dynamics that converge to the copies-core solution. Since we are aiming to reach the copies-core solution, we work with the expanded graph as defined above. The following definitions will be used in the subsequent algorithms. Furthermore, we will overload in the dynamics to be the agentsβ allocation aspirations for the copies. In our dynamics, an edge is formed only if these aspirations can be fulfilled by such match.
@@
The following is an assumption that we are going to consider for both the centralized and distributed algorithms.
@@
If a given -matching problem does not satisfy this assumption, the degree of the violating nodes still can not exceed the number of nodes in the opposite class in any -matching. Hence, the problem remains the same if we decrease the -values of these nodes just enough to satisfy the above assumption.
The algorithms use the aforementioned functions, however, the superscript is suppressed and considered as the matching at the current iteration.
@@ III PATHS TRANSFERS ALGORITHM
Initialize the matching to be empty and the aspirations of all copies to be . Set to be the largest discretization value for all the weights of the edges.
@@
Augmenting path case: If there exists a directed path in from the considered copy to some copy then let and disregarding the directions of the edges. Add the edges in to and remove the edges in from .
This case will lead to decreasing by or depending on whether is greater than or equal to zero.
@@
Copies exchange case: If there exists a directed path in from the considered copy to some copy with then let and disregarding the directions of the edges. Add the edges in to and remove the edges in from
This case will lead to decreasing by 1 as a copy with non-zero aspiration will be matched instead of a copy with zero aspiration.
@@
The inner loop of step must end in a finite time. The inner loop will end if it goes through the augmenting path or copies exchange cases or if the considered copy reaches zero aspiration in the decreasing aspiration case. The inner loop cannot infinitely go through the decreasing aspiration and the aspiration transfer case since the total aspirations of copies in strictly decreases with each implementation of these cases and it is bounded below by zero. Hence, in a finite number of steps, will decrease by either the loop going through the augmenting path or copies exchange cases or the considered copy reaching zero aspiration.
@@
The outer loop of step ends in a finite time because strictly decreases with each iteration.
@@
Note that the Paths Transfers algorithm is polynomial in the input parameters including the weight values but can be modified to be polynomial in and and exclude the weight values dependency. To achieve that, the aspiration transfer case should be modified such that the copies aspirations change by the minimum change of aspiration that leads to a new edge in the equality graph to appear or for one of the copies to hit zero aspiration. The decreasing aspiration case should be modified such that the decrease in aspiration is the minimum decrease that either leads the reconstructed to contain an edge from the considered copy to any other copy or leads to reach zero.
@@
In this section, we introduce the -Matching Proposals dynamics. Intuitively, the algorithm follows these steps:
@@
Agents come with arbitrary non-negative initial aspirations for all their copies. The aspirations value should lie on a grid of width , a discretization value for all the weights of the edges.
@@
If the current agents are matched, the proposal is skipped, otherwise, the proposing agent asks for and receives an aspiration value from the other agent.
@@
The receiving agent sends the minimum aspiration of its unmatched copies if any exist, otherwise, it sends the minimum aspiration of the matched copies. Similarly, the proposing agent considers the minimum aspiration of its unmatched copies if any exist, otherwise, it considers the minimum aspiration of the matched copies. If the total of the received aspiration in addition to the proposing agentβs considered aspiration raised by is less than or equal to the proposed match weight, a match between the two copies of the considered aspirations is formed, otherwise, the proposal fails.
@@
If a match is successfully formed, the proposing agent increases its aspiration to be the difference between the weight of the edge joining the proposer and the receiver and the receiverβs aspiration. If any of the copies in the new match were already matched, they dissolve their previous matches first to join the new one.
@@
The proof follows three steps illustrated in the following subsections. First, starting from any feasible state, there exists a sequence of proposals that reaches a copies-core solution within some finite time . Second, the probability of the occurrence of such sequence goes to one. Third, a state is absorbing if and only if it is a copies-core solution.
I. There exists a sequence of proposals that reaches a copies-core solution from any feasible state
@@ II Borel-Cantelli lemma
The sequences in propositions and occur with a positive probability, bounded away from zero, within some finite time that does not depend on the state. Hence, by the Borel-Cantelli lemma, the probability of the occurrence of a sequence of proposals that lead to the copies-core solution goes to one as time goes to infinity.
III. A state is absorbing in the -Matching Proposals dynamics if and only if it is a copies-core solution
@@
and the zero-gain for unmatched copies condition is satisfied. Otherwise, such that and due to Assumption there exists a node in the opposite class that is not connected to through any copy. If proposes to match with such node, the proposal must fail since the state is absorbing and the matching structure cannot change, yet that leads to one copy from decreasing its aspiration, which changes the state, and hence provides a contradiction.
@@
Given , then and the stability against pairwise deviation is satisfied. Otherwise, and such that Thus, if proposes to or vice versa, the proposal must succeed, consequently changing the matching and the aspirations, which establishes that the state is not absorbing and provides a contradiction.
@@
Assume that the state is a copies-core solution at iteration . Assume the proposal happening at iteration is from some node to any node
@@
If and were already connected, no change occurs.
@@
If and are not connected, then by the stability against pairwise deviation property, and Hence, all proposals must fail and no change can occur in the matching. In addition, from the zero-gain for unmatched copies condition, , then none of the copies can decrease its aspiration.
Similar justification applies if a node in proposes to any node in Hence, is absorbing.
@@ VI ILLUSTRATION
In this section, we demonstrate the distributed learning dynamics on the aforementioned multi-agent task assignment setting as follows.
@@
We consider a set of tasks, where
@@
each task can afford to utilize up to robots.
@@
We consider a set of robots, where
@@
each robot has accuracy such that and
@@
each robot can perform up to tasks.
@@
In our setup, we are assuming a negative correlation with the sophistication of the robot and how many tasks it can perform
@@
The value of a matching of task with robot is where is proportional to the accuracy of the robot multiplied by the value of the task.
We applied the -Matching Proposals dynamics and examined the evolution of the matching and copiesβ aspirations.